# Goldilocks Isolation: High Performance VMs with Edera

### v1

@@ 6 Analysis
6.1 Performance
We compare the performance of Edera under both to several existing systems. Docker using runc acts as a baseline metric for the performance of operating system virtualization. gVisor and Kata Containers (Kata) allow a comparison with other systems with strong isolation. We especially compare Edera in PV mode to Docker and gVisor, as none of these systems require virtualization extensions, while comparing Edera PVH with Kata to compare performance when such extensions are available. We thus run all systems on an OVH virtual machine with virtualization extensions, and run Docker and gVisor benchmarks both on bare metal and on the Edera hypervisor to separate out the performance impact of the Edera hypervisor from that of running benchmarks in Edera zones.
To assess performance, we evaluate both startup and runtime. Within runtime performance, we look at the CPU and memory usage, as well as the speed of various system calls. To estimate real-world performance, we also evaluate the runtime of each system for some example workloads. We run all tests 5 times and present the average of these runs, with the standard error shown in all graphs.
Figure 5: A comparison of the memory speed in MB/s using metrics from both sysbench and perf-bench. Higher is better.
Figure 6: A comparison of the CPU speed in events per second using metrics from sysbench. Higher is better.
Startup time. We first examine the startup time of each system, shown in table 2. Edera takes 0.77 seconds to startup in PV mode, compared to 0.97 seconds in PVH mode. This is a slower startup time than Docker and gVisor, which start up in 0.18 and 0.2 seconds respectively, but faster than Kata Containers, which starts up in 1.93 seconds. However, Edera zones can be started before they are needed and kept running in an idle state until use. In this way, the user does not need to wait for the 0.77 second startup time each time they use a zone. These idle warm zones can initially be allocated no memory or CPU resources, then resources can be dynamically allocated once the zone becomes active due to the flexibility of the Edera hypervisor.
@@ 6.1 Performance
Docker 177.4
Docker (bare metal) 203.8
gVisor 281.8
Edera-PV 765.8
Kata Containers 1934.2
Edera-PVH 968.6
Table 2: The startup time for each system.
Memory and CPU benchmarks. Figure 5 shows the memory usage from sysbench, a benchmarking tool that performs reads and writes to a data buffer, and the speed of the memset and memcpy operations from the perf-bench utility. Edera outperforms the other systems on both sysbench tests, although it is only slightly faster than Docker and Kata on bare metal (.03% and 1.8% respectively). Edera also equals or outperforms Docker, gVisor, and Kata for both perf-bench operations. Note that the hypervisor does not make a large difference for memory operations, with Docker performing similarly both on bare metal and on the Edera hypervisor.
Figure 6 illustrates the performance of the sysbench CPU benchmark. For this test sysbench monitors CPU usage while checking prime numbers up to 150,000 over 4 threads. Edera is slightly slower than both Docker and Kata on bare metal (.9% and .7% slower respectively). Docker runs much slower on the Edera hypervisor than on bare metal, illustrating that some of the CPU performance lost through the use of vCPUs in the Edera hypervisor is made up through the CPU management by Edera zones. Edera benefits from using paravirtualization for resource management, as it allows for faster memory mapping and CPU scheduling by using a software interface instead of full hardware emulation. Edera has a full view of the hardware, allowing smart optimizations like co-locating pages in a zone.
@@
Figure 7: A comparison of the operations per second of system calls. Note that this graph is shown on a logarithmic scale. Higher is better.
@@
Figure 8: A comparison of the system call latency. Note that this graph is shown on a logarithmic scale. Lower is better.
Figure 9: A comparison of runtime of kcbench in milliseconds. Lower is better.
@@
Figure 10: A comparison of runtime of nginx in milliseconds. Note that this graph is shown on a logarithmic scale. Lower is better.
@@
System calls. Next, we evaluate the speed of executing several popular system calls. The mechanism by which system calls are passed to the host system differs widely between the systems we evaluate, so we determine how these approaches impact performance. Figure 7 shows the operations per second of each system running the fork(), execve(), getpgid(), and epoll() system calls based on results from perf bench. This graph shows that Edera outperforms both Docker and gVisor for the execve and epoll operations, and outperforms gVisor for fork and getpgid. Edera avoids much of the indirection used by gVisor, leading to much faster performance. Kata outperforms Edera for fork(), execve(), and getpgid(). However, in PVH mode, Edera performs fork() and execve() similarly to Kata. Both Kata and Edera PVH use virtualization extensions to achieve this higher system call performance. Figure 8 further shows the latency of fork(), execve(), and getpgid(). The latency of Edera in PVH mode is similar to that of Docker and Kata. Docker also has lower latency when run under the Edera hypervisor, illustrating the scheduling performance of the hypervisor.
@@
Real-world workloads. Finally, we evaluate the runtime of workloads in each system to understand how they perform in real-world environments. These runtimes are shown in fig. 9 and fig. 10. We first discuss the time to build a kernel in each system using kcbench. Edera’s kernel build time is 5% slower than Docker, 59% faster than gVisor, and 52% faster than Kata. This kernel build benchmark represents a more realistic workload with several different system calls and operations. Next, we evaluate the runtime of a simple nginx server using Apache Bench. We find that Edera outperforms gVisor by a factor of 10 (18 ms compared to 220 ms for Docker), and performs comparably to Docker (15 ms). Edera-PVH has slightly faster performance than Kata according to this metric, although Edera-PV is slightly slower. The use of a webserver like nginx is a common application for cloud applications.
6.1.1 Summary
@@
In summary, Edera outperforms gVisor in all performance metrics tested with the exception of startup time, while overall performing only slightly worse than Docker and similarly to Kata. This means that Edera achieves strong application isolation with minimal impact on performance even for systems that do not have access to virtualization extensions. The system startup time performs the worst, with 0.77 seconds of startup time for Edera compared to 0.2 seconds for Docker. When virtualization extensions are available, Edera in PVH mode has improvements to system call runtime over PV mode, and outperforms Kata on memory and CPU operations.
6.2 Security
Attack Mitigation Protected?
CVE-2022-0185 No shared kernel
CVE-2022-0492 No shared kernel
CVE-2022-0811 (cr8escape) No shared kernel
CVE-2022-0847 (Dirty Pipe) No shared kernel
CVE-2022-23648 No shared kernel
CVE-2024-0132 Driver isolation
CVE-2024-21626 (Leaky Vessels) Runtime exploits can’t reach host
@@ 6.2 Security
Table 3: Overview of recent container escape CVEs and how they are mitigated by Edera.
Next, we evaluate the security of Edera using our requirements from section 3 and previous container escape attacks.
6.2.1 Requirements
Edera protect meets the requirements from section 3:
@@ 6.2.1 Requirements
  • Eliminate shared kernel. Edera uses a type-1 hypervisor to isolate containers into zones, with each zone containing its own guest operating system. This separation eliminates the shared kernel used by OS virtualization.
@@
  • Use existing images. Edera is compatible with existing OCI images, and protect-cri complies with the Kubernetes CRI specification for seamless use by existing systems. This reduces the barriers to use by ensuring that existing applications, even those using proprietary software, can be run in Edera without modification.
@@
  • No specialized hardware required. Edera works on all commercially available hardware, even that without virtualization extensions. It does so by using paravirtualization on guest kernels to allow full use of Linux system calls without the need for hardware virtualization.
  • Driver isolation. Edera allows hardware drivers to run in isolated zones, protecting the hypervisor in the event of a driver compromise.
  • Runtime performance. Edera achieves faster runtime performance than gVisor by all metrics and performs within 15% by all metrics when compared to Docker with the exception of getgpid. This meets the requirement specified in section 3. When run in PVH mode, Edera achieves even better system call and workload performance.
  • Startup performance. Edera adds 0.56 seconds of startup time compared to our baseline of Docker with runc. Although this startup time does not meet the requirement from section 3, it can be further reduced by using warm zones to meet this requirement.
6.2.2 Preventing container escapes
@@
Next, we evaluate Edera’s security by examining how it can be used to prevent container escapes. We evaluate seven container escapes from 2022-2024 to determine how Edera fares against these attacks. We focus on vulnerabilities in Linux in this timeframe, as these could impact both Edera and other isolation techniques and leave out analysis of vulnerabilities in other isolation techniques. Table 3 shows the results.
@@
Findings. We find that Edera prevents all of these attacks through a combination of having no shared kernel and driver isolation. The prevention mechanism is labeled in table 3 for each attack. Without a shared kernel, attacks on capabilities, namespaces, and other OS isolation techniques do not give an attacker access to the host machine. This means that any kernel exploit that previously could cause a container escape will only impact the guest OS on Edera. As such, Edera is not vulnerable to any of these past container escape attacks and is further not vulnerable to this class of attack in the future. We provide in-depth explanations of three of these CVEs to illustrate how Edera prevents them.
@@
CVE-2024-21626. A vulnerability in runc, known as “Leaky Vessels,” uses a leak of an internal file descriptor to create a container with a working directory in the host filesystem namespace. This can be used to gain access to the host filesystem, or create a privilege escalation by re-writing files on the host filesystem. In Edera, guest OSs do not share a filesystem with the hypervisor, so even if a guest OS had access to a hypervisor file descriptor, they would not be able to access or overwrite files on the hypervisor.
@@ 7.1 Zone Flexibility
7.2 Usability
The good performance of Edera will allow organizations to use security mechanisms for all containers, not just those related to security. For performance reasons, many organizations use OS virtualization for most operations, and a system with stronger isolation (like gVisor or Kata Container) for certain, security critical applications. This segmentation of operations leaves most workloads vulnerable to container escapes. Edera obviates the need for this segmentation, providing strong security for all applications.
Edera achieves this by providing a drop-in replacement for the Kubernetes runtime so that it can be used for isolation without sacrificing usability at all other layers of Kubernetes. To demonstrate, consider a simple Kubernetes configuration:
apiVersion: v1
kind: Pod
metadata:
  name: leaky-vessel
  namespace: test
spec:
  containers:
  - name: leaky-vessel
    image: example/leaky-vessel:0.1
    imagePullPolicy: Always
    env:
    - name: SUPER_ORCHESTRATOR_SECRET
      value: "this-is-fine-leaky"
By adding a 2 line annotation to this configuration, the application is now protected by Edera (duplicate configuration is removed for brevity):
metadata:
  annotations:
    dev.edera/kernel: \
    "ghcr.io/edera-dev/linux-kernel:latest"
    dev.edera/memory: "600"
@@ 7.3 Future work
  • Further optimizations to improve Edera’s startup and runtime performance.
  • Expanding Edera’s driver isolation to non-NVDIA GPUs. Other GPUs have different supported drivers, and do not support vGPU used by our implementation.
@@
  • While Edera limits kernel exploits to a single zone, it could be expanded to provide protection within the zone through the use of monitoring. A protocol like expanded Berkeley Packet Filtering (eBPF) could be used in the hypervisor to detect anomalies and prevent exploits running in a zone. Further, monitoring of known compromised kernels could be used for compromise detection.
8 Conclusion
We present and evaluate Edera, a type-1 hypervisor that has performance comparable to popular operating system virtualization technologies while providing strong isolation. We implement Edera for both a Kubernetes compatible container runtime, and for driver isolation to demonstrate how it can work within existing infrastructure. We find that Edera outperforms gVisor in CPU performance, memory performance, and under real-world workloads, while only adding 0.48 seconds of startup time without the use of virtualization extensions. It further achieves comparable performance to Docker and Kata Containers. By eliminating the shared kernel, Edera is able to prevent all tested container escape attacks.

### v2

@@ 6. Analysis
6.1. Performance
We compare the performance of Edera to existing systems for running containers. Docker using runc acts as a baseline metric for the performance of OS virtualization. gVisor, Kata Containers (Kata), and Firecracker allow comparison with other systems with strong isolation. We run gVisor with the default configuration, which uses seccomp-based systrap platform. We run Kata Containers using the KVM hypervisor with the Dragonball VMM. We choose Dragonball as it is optimized for container workloads (Kata Containers Maintainers, 2025). See Appendix A for some performance numbers that verify this. We run Firecracker on the KVM hypervisor with the firecracker-containerd container runtime. We especially compare Edera in PV mode to Docker and gVisor, as none of these systems require virtualization extensions, while comparing Edera PVH with Kata and Firecracker to compare performance when such extensions are available. We run all experiments on a bare metal OVH machine with virtualization extensions. The machine has 16 GB of memory and an Intel Xeon CPU with 8 cores at 3.20GHz. There were 3 boot configurations we used to run these benchmarks: kvm, edera-pv, and edera-pvh. We ran benchmarks on all systems where they would run. Docker and gVisor run in userspace and are not hypervisor or operating system dependant, so these benchmarks ran in all three boot configurations. Edera, Firecracker, and Kata depend on the hypervisor, so were only run on the appropriate systems.
We evaluate both startup and runtime performance. Within runtime performance, we look at the CPU and memory usage, as well as the speed of various system calls. To estimate real-world performance, we also evaluate the overall runtime for some example workloads. All benchmarks are run on OCI container images (generally the latest Alpine image), with a fresh container created for each run to model the performance in a real-world cloud application. Benchmarks are run with access to 4 vCPUs and 4096 MB of memory, except for the kcbench and pgbench which required more resources to run successfully on all systems. We run benchmarks 5 times and present the average of these runs, with the standard error shown in all graphs.
@@ 6.1. Performance
Docker (pv) 381
Docker (pvh) 378.8
gVisor (pvh) 444
gVisor (pv) 452.2
Docker (kvm) 571.6
gVisor (kvm) 684.4
Edera-PV 1120.8
Kata Containers 1913
Firecracker 2421.8
Edera-PVH 2516.8
Table 1. The startup time for each system. Rows are sorted by time.
Figure 3. A comparison of the memory speed in GB/s using metrics from both sysbench and perf-bench. Higher is better.
system Speed (events/sec)
Firecracker 183.8
Docker (pvh) 276.62
gVisor (pvh) 310.84
gVisor (pv) 328.11
Docker (pv) 329.03
Edera-PVH 359.41
Edera-PV 359.95
Kata Containers 366.67
gVisor (kvm) 366.89
Docker (kvm) 367.84
Table 2. A comparison of the CPU speed in events per second using metrics from sysbench. Higher is better, results are sorted.
Startup time. We first examine the container startup time, including the full time from VM init to starting the container, shown in table 1. Container images were cached, so this does not include downlaod time. Edera takes 1.1 seconds to startup in PV mode, compared to 2.5 seconds in PVH mode. This is a slower startup time than Docker and gVisor on Linux, which start up in 0.57 and 0.68 seconds respectively, but PV is faster than Kata Containers and Firecracker, which start up in 1.9 and 2.4 seconds respectively. Note that this is slower than the Firecracker startup time reported in (Agache et al., 2020) due to our reporting the complete container startup time rather than just the VM boot time.
However, Edera zones, as well as Kata or Firecracker VMs, can be started before they are needed and kept running in an idle state until use. In this way, the user does not need to wait for the 1 second startup time to use a zone. With Edera’s dynamic resource allocation these idle warm zones can initially be allocated no memory or CPU resources, then resources can be dynamically allocated once the zone becomes active.
Memory and CPU benchmarks. We assess the memory and CPU speed of each system. Each system was configured with the same resources, so this helps compare their efficiency. Figure 3 shows the results from sysbench, a benchmarking tool that performs reads and writes to a data buffer, and the speed of memset and memcpy from perf-bench. sysbench used a block size of 1 GB and total memory size of 20 GB for memory operations and sequential write. For the perf-bench memory benchmarks, the size was set to 100MB.
Edera outperforms the other systems on both sysbench tests, although by a small margin. Edera also demonstrates the best performance for memcpy and memset, with memcpy showing noticably slower results for Firecracker and gVisor. Docker performs similarly on bare metal and on the Edera hypervisor indicating minimal performance impact from the hypervisor on memory operations. gVisor performs slightly worse on the Edera hypervisor due to the additional indirection. In summary, all the systems, with the exception of gVisor and Firecracker have very similar memory performance.
Table 2 illustrates the performance of the sysbench CPU benchmark. For this test sysbench monitors CPU usage while checking prime numbers up to 150,000 over 4 threads. Edera-pv is 2.14% slower than Docker on Linux. Docker is 10-25% slower on the Edera hypervisor, illustrating some performance loss through the use of vCPUs in Edera. Edera itself was faster on the hypervisor, illustrating that some of this performance is made up through the CPU management by Edera zones. Firecracker performed at roughly half the speed, likely due to inefficient use of the configured 4 vCPUs. More fine-grained CPU configuration to improve this is not supported in firecracker-containerd.
@@
Figure 4. A comparison of the operations per second of system calls shown on a logarithmic scale. Higher is better.
@@
Figure 5. A comparison of the system call latency. Note that this graph is shown on a logarithmic scale. Lower is better.
Figure 6. A comparison of runtime of kcbench in milliseconds. Lower is better.
@@
Figure 7. A logarithmic scale comparison of runtime of nginx in milliseconds. Lower is better.
@@
System calls. Next, we evaluate the speed of executing several popular system calls. The mechanism by which system calls are passed to the host system differs widely between the systems we evaluate, so we determine how these approaches impact performance. Figure 4 shows the operations per second of each system running fork(), execve(), and getpgid() based on results from perf bench. This graph shows that Edera outperforms gVisor in all system calls, and performs slower than Docker on bare metal (for example fork() is 41.8% slower on Edera than Docker). Edera avoids much of the indirection used by gVisor, leading to much faster performance. However, with virtualization extensions, Edera PVH is able to outperform Docker (with fork() 169.4% faster than Docker). This is slower than the system call performance of Kata and Firecracker, which also use virtualization extensions, with Kata 34.4% faster and Firecracker 43.8% faster than Edera PVH.
Figure 5 further shows the latency of fork(), execve(), and getpgid(). Similar to the system call performance, the latency of Edera PVH is better than Docker and gVisor, but worse than Kata and Firecracker.
@@
Real-world workloads. Finally, we evaluate the runtime of workloads in each system to understand how they perform in real-world environments. These runtimes are shown in fig. 6 and fig. 7. We first evaluate kernel build time using kcbench. We set the kernel version to 6.11.3 and 8 jobs in parallel. For this benchmark, all systems had access to 8 vCPUs and 16384 MB of memory. Edera’s kernel build time is 5% slower than Docker, 59% faster than gVisor, and 52% faster than Kata. A kernel build represents a more realistic workload with several different system calls and operations. As this benchmark requires high memory usage, it does not run on firecracker-containerd. Next, we evaluate the runtime of a simple nginx server using Apache Bench over 10,000 requests. We find that Edera outperforms gVisor by a factor of 10 (18 ms compared to 220 ms for gVisor), and performs comparably to Docker (15 ms). Edera-PVH has slightly faster performance than Kata according to this metric, although Edera-PV is slightly slower. The use of a webserver like nginx is a common application for cloud applications.
6.1.1. Summary
@@
In summary, Edera outperforms gVisor in all performance metrics with the exception of startup time, while overall performing only slightly worse than Docker and similarly to Kata and Firecracker. This means that Edera achieves strong application isolation with minimal impact on performance even for systems that do not have access to virtualization extensions. The system startup time performs the worst, with 1.1 seconds of startup time for Edera compared to 0.38 seconds for Docker. Overall memory and CPU performance for Edera is similar to that of Docker, and system calls perform worse. When virtualization extensions are available, Edera in PVH mode has improvements to system call runtime over PV mode, and outperforms Kata and Firecracker on memory and CPU operations.
6.2. Requirements
Attack Protected?
CVE-2022-0185
CVE-2022-0492
CVE-2022-0811 (cr8escape)
CVE-2022-0847 (Dirty Pipe)
CVE-2022-23648
CVE-2024-0132
CVE-2024-21626 (Leaky Vessels)
@@ 6.2. Requirements
Table 3. Overview of recent container escape CVEs and how they are mitigated by Edera.
Next, we evaluate Edera using our requirements from section 3.
@@
Eliminate shared kernel. Edera uses a type-1 hypervisor to use CPU partitioning to isolate containers in zones that have their own guest OS. This separation eliminates the shared kernel used by OS virtualization.
@@
Use existing OCI images. Edera is compatible with OCI images, and protect-cri complies with the Kubernetes CRI specification. This ensures that existing applications, even those using proprietary software, can be run in Edera without modification.
@@
Cloud compatibility. Edera does not rely on direct access to virtualization extensions. It uses paravirtualization on guest kernels to allow full use of Linux system calls without the need for hardware virtualization.
Memory safety. Edera is written in Rust and uses MISRA C components from Xen, providing memory safety.
Runtime performance. Edera performs within 5% of Docker by CPU and memory benchmarks. Some system calls were above this threshold, but when run in PVH mode Edera is faster than Docker by these metrics as well.
Startup performance. Edera adds 0.74 seconds of startup time compared to Docker, meeting the requirement of less than 1 second of additional overhead. Edera’s dynamic resource allocation means it can be further reduced with warm zones that are given resources when needed.
6.3. Security
@@
Next, we evaluate Edera’s security by examining if it can prevent container escapes. We evaluate seven container escapes from 2022-2024 to determine how Edera fares against these attacks. We focus on vulnerabilities in Linux, as these could impact both Edera and other isolation techniques and omit analysis of vulnerabilities specific to other isolation techniques. Table 3 summarizes the results.
@@
Findings. We find that Edera prevents all attacks through the use of hypervisor virtualization. Without a shared kernel, attacks on capabilities, namespaces, and other OS isolation techniques do not give an attacker access to the host machine. This means that any kernel exploit that causes a container escape will only impact the guest OS on Edera. As such, Edera is not vulnerable to any of these past container escape attacks and is further not vulnerable to this class of attack in the future. We provide in-depth explanations of three of these CVEs to illustrate how Edera prevents them.
@@
CVE-2024-21626. A vulnerability in runc, known as “Leaky Vessels,” uses a leak of an internal file descriptor to create a container with a working directory in the host filesystem namespace. This can be used to gain access to the host filesystem or create a privilege escalation by re-writing files on the host filesystem. In Edera, guest OSs do not share a filesystem with the hypervisor, so even if a guest OS had access to a hypervisor file descriptor, they would not be able to access or overwrite files on the hypervisor.
@@ 7.1. Zone Flexibility
7.2. Usability
The good performance of Edera will allow organizations to use security mechanisms for all containers, not just those related to security. For performance reasons, many organizations use OS virtualization for most operations, and a system with stronger isolation (like gVisor or Kata Containers) for certain, security critical applications. This segmentation of operations leaves most workloads vulnerable to container escapes. Edera obviates the need for this segmentation, providing strong security for all applications.
Edera achieves this by providing a drop-in replacement for the Kubernetes runtime so that it can be used for isolation without sacrificing usability at all other layers of Kubernetes. Edera can be applied to a Kubernetes configuration with a 1 line change adding the runtime class to the spec section:
@@ 7.3. Future work
  • Further optimizations to improve Edera’s startup and system call performance.
  • Using Edera for driver isolation. Isolating GPUs and other drivers further reduces the TCB.
@@
  • A protocol like expanded Berkeley Packet Filtering (eBPF) could be used in the hypervisor to detect anomalies and prevent exploits running in a zone. Further, monitoring of known compromised kernels could be used for compromise detection.
8. Conclusion
We present and evaluate Edera, an optimized VMM that has performance comparable to popular operating system virtualization technologies while providing strong isolation. We implement Edera along with a Kubernetes compatible container runtime to demonstrate how it can work within existing infrastructure. We find that Edera outperforms gVisor in CPU performance, memory performance, and under real-world workloads, while only adding 0.74 seconds of startup time without the use of virtualization extensions. It further achieves comparable performance to Docker, Kata Containers, and Firecracker.
Appendix A Kata Backend Comparison
In order to select a Kata backend for the comparisons in this paper, we confirmed which had the best performance in our setup. We tested clh, cloud hypervisor, dragonball, and qemu. We found the results to be very similar, with Dragonball performing slightly better in a few cases. The CPU, memory, and syscall performance from this comparison are shown in fig. 8, fig. 9, and fig. 10.
Figure 8. A comparison of the CPU speed in events per second using metrics from sysbench. Higher is better.
Figure 9. A comparison of the memory speed in GB/s using metrics from both sysbench and perf-bench. Higher is better.
Figure 10. A comparison of the operations per second of system calls shown on a logarithmic scale. Higher is better.